xdy + (y-x^2y^2)dx = 0


1
Expert's answer
2021-07-16T03:27:27-0400
"y'+\dfrac{1}{x}y=xy^2"

First order Bernoulli ODE


"t=y^{1-n}=y^{1-2}=y^{-1}"

"y=t^{-1}, y'=-t^{-2}t'"

"-t^{-2}t'+\dfrac{1}{x}t^{-1}=xt^{-2}"

"-t'+\dfrac{1}{x}t=x"

"\mu=\dfrac{1}{x}"

"-\dfrac{1}{x}t'+\dfrac{1}{x^2}t=1"

"(\dfrac{1}{x}t)'=\dfrac{1}{x}t'-\dfrac{1}{x^2}"

"(\dfrac{1}{x}t)'=-1"

"\dfrac{1}{x}t=-x+c_1"

"t=-x^2+c_1x"

"y=\dfrac{1}{-x^2+c_1x}"